OpenStudy (liz_beauty45):

What is the solution to 19 = z + 12?

OpenStudy (anonymous):

Subtract 12 from both sides.

OpenStudy (anonymous):

Then you'll get your answer.

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 19=z+12 Step 1: Flip the equation. z+12=19 Step 2: Subtract 12 from both sides. z+12−12=19−12

OpenStudy (liz_beauty45):

thanksssssssssss

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 19=z+12 Step 1: Flip the equation. z+12=19 Step 2: Subtract 12 from both sides. z+12−12=19−12 z=7 Answer: z=7

OpenStudy (adelyn_helps):

http://www.geteasysolution.com/19=z+12

OpenStudy (adelyn_helps):

that website helps

rishavraj (rishavraj):

u need to isolate the variable on one side of the equality sign.. so in this question u simply need to subtract 12 on both sides

OpenStudy (jjcoolja12):

uh @rishavraj the question has already been answered

OpenStudy (liz_beauty45):

yeah i already got my answer but thanks anywayssss

OpenStudy (adelyn_helps):

its explains the answer! just so you know!

OpenStudy (jjcoolja12):

and a medal i dont get